OpenStudy (anonymous):

HELP ME PLEASE The function f(x) = −x2 + 48x − 432 models the profit, in dollars, a game developer makes for selling a new game, where x is the number of games sold and f(x) is the amount of profit. Part A: Determine the vertex. What does this calculation mean in the context of the problem? (5 points) Part B: Determine the x-intercepts. What do these values mean in the context of the problem? (5 points)

OpenStudy (anonymous):

@laughoutloud

OpenStudy (igreen):

Use \(\sf -\dfrac{b}{2a}\) to find the x-value of the vertex.

OpenStudy (igreen):

The function is in the form of \(\sf ax^2 + bx + c\).

OpenStudy (triciaal):

do you know how to complete the square to put the function in vertex form? a(x-h)^2 + k = 0 h = -b/(2a) the vertex will be (h, k)

OpenStudy (anonymous):

ok

OpenStudy (igreen):

@triciaal We don't nee it in vertex form to use \(\sf -\dfrac{b}{2a}\)

OpenStudy (igreen):

need*

OpenStudy (anonymous):

Alright .-.

OpenStudy (triciaal):

the vertex is at is the highest or lowest y-value when a is negative the parabola opens down so your vertex is at the maximum y

OpenStudy (triciaal):

@iGreen -b/(2a) is your h

OpenStudy (igreen):

\(\sf \color{red}ax^2 + \color{blue}bx + \color{lime} c\) \(\sf \color{red}{-1}x^2 + \color{blue}{48}x \color{lime}{- 432}\)

OpenStudy (igreen):

Yes, then after we find 'h', we can plug it back into \(\sf -x^2 + 48x - 432\).

OpenStudy (igreen):

To find 'k'

OpenStudy (anonymous):

OK :D

OpenStudy (igreen):

Yep, just plug in the values into \(\sf -\dfrac{b}{2a}\) first.

OpenStudy (triciaal):

@DarkMoonZ do you have any questions about my response?

OpenStudy (anonymous):

No I have none thanks guys

OpenStudy (anonymous):

Triciall give Igreen a medal for me XD

OpenStudy (igreen):

There's no need. If you want, you can tag us back to check your work.

OpenStudy (anonymous):

Ok thanks

OpenStudy (triciaal):